# Discrete Diffusion VLA: Bringing Discrete Diffusion to Action Decoding in Vision-Language-Action Policies

### v1

@@ Abstract
Vision–Language–Action (VLA) models adapt large vision–language backbones to map images and instructions to robot actions. However, prevailing VLA decoders either generate actions autoregressively in a fixed left-to-right order or attach continuous diffusion or flow matching heads outside the backbone, demanding specialized training and iterative sampling that hinder a unified, scalable architecture. We present Discrete Diffusion VLA, a single-transformer policy that models discretized action chunks with discrete diffusion and is trained with the same cross-entropy objective as the VLM backbone. The design retains diffusion’s progressive refinement paradigm while remaining natively compatible with the discrete token interface of VLMs. Our method achieves an adaptive decoding order that resolves easy action elements before harder ones and uses secondary re-masking to revisit uncertain predictions across refinement rounds, which improves consistency and enables robust error correction. This unified decoder preserves pretrained vision–language priors, supports parallel decoding, breaks the autoregressive bottleneck, and reduces the number of function evaluations. Discrete Diffusion VLA achieves 96.3% avg. SR on LIBERO, 71.2% visual matching on SimplerEnv–Fractal and 49.3% overall on SimplerEnv–Bridge, improving over both autoregressive and continuous diffusion baselines. These findings indicate that discrete-diffusion action decoder supports precise action modeling and consistent training, laying groundwork for scaling VLA to larger models and datasets.
@@ 1 Introduction
Vision–Language–Action (VLA) models enable robots to interpret visual and linguistic inputs and execute corresponding action sequences. Modern VLA frameworks typically adapt a large pretrained vision–language model (VLM) by adding an action-generation head that outputs motor commands (either continuous trajectories or discrete tokens). Current approaches fall into two paradigms: (1) an autoregressive (AR) approach, inspired by GPT-style transformers, that predicts action tokens sequentially (e.g., OpenVLA (Kim et al., 2024), $\pi_{0}$-FAST (Pertsch et al., 2025)), and (2) a continuous diffusion approach that treats an entire action trajectory as a continuous signal to iteratively denoise (e.g., $\pi_{0}$ (Black et al., 2024) and SmolVLA (Shukor et al., 2025)). Continuous diffusion can model sophisticated multimodal actions better than AR but remains decoupled from the VLM backbone. Even partial integration efforts (e.g., the Transfusion (Zhou et al., 2024) design in $\pi_{0}$) still rely on diffusion-specific training and iterative sampling, lacking a truly unified structure which is consistent with VLM part.
@@
Drawing on recent advances of discrete diffusion and discrete flow‑matching for language and multi-modal generation (Nie et al., 2025a; Shi et al., 2024b; Gat et al., 2024; Kim et al., 2025a; Yang et al., 2025), we introduce Discrete Diffusion VLA, the first VLA framework to unify vision, language, and action in a single transformer, maintaining strong action-modeling capability. In Discrete Diffusion VLA, each action dimension is discretized into tokens via a binning scheme (as in RT-1 (Brohan et al., 2022) and OpenVLA (Kim et al., 2024)) and grouped into fixed-length chunks. The fixed-length action tokens are exactly suitable for discrete diffusion models. During training, we mask a subset of these action tokens and train the transformer to predict them from the context of the unmasked tokens across all modalities. The noise scheduling is governed by intuitive mask ratios. At inference, all action tokens start masked and are iteratively decoded by repeatedly predicting tokens and re-masking low-confidence ones until convergence, following the “first‑easy, then‑hard” philosophy. Moreover, we employ a second re-masking technique to guarantee the consistency across different denoising steps. This yields flexible parallel decoding and robust error correction.
@@
For robotics manipulation tasks, in contrast to continuous diffusion decoders, our formulation keeps action generation inside a unified transformer with the same training objective (i.e., cross-entropy loss) used by VLMs. This preserves much of the backbone’s pretrained vision and language capability, analogous to extending a vocabulary to new languages, while offering a potential path to inherit from unified transformer’s scaling behavior, paving the way for large‑scale VLA research. Moreover, Discrete Diffusion VLA breaks the AR decoders’ left-to-right bottleneck. Action chunks are adaptively decoded in parallel over a small, fixed number of steps, and unconfident tokens can be revisited via iterative re-masking, leveraging full cross-modal context (including inter-action dependencies) for refinement.
@@
We evaluate Discrete Diffusion VLA across three different robot settings: a Franka Panda arm on LIBERO (Liu et al., 2023), a Google Robot in SimplerEnv–Fractal (Li et al., 2025), and a WidowX Robot in SimplerEnv–Bridge (Li et al., 2025). Using only RGB inputs, language and end-effector positions (no depth or affordances), Discrete Diffusion VLA achieves 96.3% average success rate on LIBERO ( +0.9% vs. OpenVLA-OFT (Discrete) ), 71.2% visual matching and 64.1% overall on SimplerEnv–Fractal, and 49.3% overall on SimplerEnv–Bridge ( $\approx$  +9.8% vs. $\pi_{0}\text{and}\pi_{0}$+FAST). Our model outperforms both diffusion and AR baselines on most tasks while using fewer Number of Function Evaluations (NFEs) than autoregressive paradigm. Ablations confirm the benefits of our adaptive decoding strategy.
@@
In summary, our contributions are threefold: 1) We introduce the first discrete diffusion action head for VLA, unifying action generation with vision–language in one transformer while maintaining strong performance. 2) We develop a “first‑easy, then‑hard” adaptive decoding strategy with iterative re-masking, enabling parallel action-token decoding and error correction, improving accuracy with the unified architecture. 3) We validate Discrete Diffusion VLA on Franka Panda in LIBERO, Google Robot and WidowX Arm in SimplerEnv, obtaining 96.3% avg. SR on LIBERO, 64.1% and 49.3% overall on SimplerEnv–Fractal and –Bridge, consistently outperforming AR and continuous diffusion baselines (e.g., $\pi_{0}$).
@@ 2 Related Works
2.1 Autoregressive Vision-Language-Action Models
When the vision–language–action model was first proposed in RT-1 (Brohan et al., 2022) and RT-2 (Zitkovich et al., 2023), it followed an autoregressive (AR) paradigm that discretizes robot actions into token sequences and generates them via next-token prediction. OpenVLA (Kim et al., 2024) advanced this direction by adopting a 7B-parameter Llama-2 (Touvron et al., 2023b) backbone and fusing DINOv2 (Oquab et al., 2024a) and SigLIP (Zhai et al., 2023) visual features, demonstrating strong general-purpose manipulation. LAPA (Ye et al., 2024) reduces reliance on precise action labels by learning a VQ-VAE–based latent action space (Van Den Oord et al., 2017) from Internet videos, pretraining in this space and finetuning with a small action-labeled set, leading to improved language-conditioned control. $\pi_{0}$-FAST (Pertsch et al., 2025) introduces the FAST tokenizer, combining the Discrete Cosine Transform (DCT) (Khayam, 2003) with Byte-Pair Encoding (BPE) (Shibata et al., 1999) to compactly encode 5050Hz control signals for efficient high-frequency AR training. More recently, OpenVLA-OFT (Kim et al., 2025b) proposed parallel decoding and action chunking to rapidly adapt OpenVLA (Kim et al., 2024), while SpatialVLA augments VLA with depth cues to strengthen spatial reasoning.
In contrast to AR’s left-to-right factorization, our approach decodes action chunks in parallel and iteratively re-masks low-confidence tokens, leveraging full multi-modal and inter-action context for robust refinement across a small, fixed number of inference steps.
2.2 Diffusion-based Vision-Language-Action Models
To capture the multi-modal structure of manipulation trajectories and enable fine-grained control, recent VLA systems (Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025a; b; Wen et al., 2025; Li et al., 2024a; Liu et al., 2025; 2024) augment vision–language backbones with diffusion or flow-matching action generators (Janner et al., 2022; Chi et al., 2023; Liang et al., 2023; Ni et al., 2023; Liang et al., 2025; Chen et al., 2025a) while retaining autoregressive conditioning over vision–language tokens. These generators natively produce actions in continuous spaces. CogACT (Li et al., 2024a) introduces a lightweight diffusion-based action transformer, while RDT-1B (Liu et al., 2024) scales to 1.2B parameters and pretraining over 46 multi-robot datasets, achieving state-of-the-art bimanual dexterity. AgiBot World Colosseo (Bu et al., 2025a), UniVLA Bu et al. (2025b) and related works (Wen et al., 2025; Zhong et al., 2025; Liang et al., 2024) explore hierarchical latent actions and hybrid AR–diffusion schemes. $\pi_{0}$ (Black et al., 2024) employs flow matching to learn dense continuous trajectories and is the state-of-the-art scheme, while $\pi_{0.5}$ (Intelligence et al., 2025) adds a two-stage regime with a high-level “thinking” phase for semantic sub-task planning, enabling robust real-world generalization.
Unlike these designs that attach a diffusion/flow module alongside the VLM and rely on iterative denoising loops, our method performs discrete diffusion entirely within a single transformer over tokenized action chunks, preserving the VLM’s emergent language and multimodal capabilities and offering a potential path for VLA to scale.
@@ 2.3 Discrete Diffusion Models
Discrete diffusion has recently achieved strong results on discrete data, notably tokenized images and natural language. Foundational work D3PM (Austin et al., 2021) formalizes discrete diffusion as a Markov chain. The forward process progressively corrupts an input $\mathbf{x}_{0}$ over $T$ timesteps, where the transition to a noised state $\mathbf{x}_{t}$ is defined by a product of transition matrices $\mathbf{\bar{Q}}_{t}=\mathbf{Q}_{t}\dots\mathbf{Q}_{1}$. Different from continuous diffusion models, it factorizes across positions, so each token $\mathbf{x}_{i}$ (represented as a one-hot vector) is independently transformed into a categorical distribution that,
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\prod_{i=1}^{L}\text{Categorical}(\mathbf{x}_{t,i}|\mathbf{\bar{Q}}_{t}\mathbf{x}_{0,i}).\vskip-6.0pt$ (1)
[…]
Our work extends this line to the action modality: we perform discrete diffusion within a single transformer over tokenized action chunks, preserving language capabilities and VLM synergy. This yields competitive state-of-the-art VLA performance in our evaluations and lays groundwork toward unified, scalable, multi-modal foundation models spanning vision, language, and action.
@@ 3.1 Overview
Discrete Diffusion VLA is a single-transformer vision–language–action (VLA) policy that predicts a fixed-length future action chunk from multimodal context. Given image observations (single or multi-view) and a language instruction, we discretize each continuous control dimension into tokens and assemble them into a fixed-length action sequence. A single transformer processes all modalities: vision tokens from a frozen visual encoder, language tokens from a pretrained LM, and the action tokens. We pose action generation as masked-token denoising inside the same transformer that encodes vision and language.
During training, at each step, we sample a mask ratio over action positions, replace the selected tokens with mask token [MASK], and optimize a cross-entropy objective to recover the masked tokens, aligning action learning with discrete-diffusion style corruption and denoising. During inference, we initialize all action positions as [MASK] and run a small number of parallel refinement iterations. An adaptive re-masking policy is employed to perform multi-step denoising and refinement, while a secondary re-masking is used to further enforce consistency across iterations.
@@ 3.2 Discrete Diffusion over Action Tokens
Let a single action chunk be a length-$L$ token sequence $\mathbf{a}_{0}=(a_{0,1},\dots,a_{0,L})$, where each $a_{0,i}\in\{1,\dots,K\}$ is obtained by binning continuous controls (position, orientation, gripper) following prior VLA discretizations (Brohan et al., 2022; Kim et al., 2024). We augment the action vocabulary with a special mask token M (i.e., [MASK]), yielding V=K+1V=K{+}1 symbols and one-hot basis $\{\mathbf{e}_{1},\dots,\mathbf{e}_{K},\mathbf{e}_{\textrm{M}}\}$.
@@
Composing transition matrices yields $\bar{\mathbf{Q}}_{t}=\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$, and the corrupted distribution at time $t$ factorizes across positions:
@@
where $L$ is the length of action chunk tokens.
The reverse (denoising) process defines conditionals $p_{\theta}(\mathbf{a}_{t-1}\mid\mathbf{a}_{t},\mathbf{c})$ under multimodal (i.e., vision and language) context $\mathbf{c}$. By Bayes’ rule, for each position $i$,
@@
where $\pi_{\theta}(i\mid\mathbf{c})$ is the model’s predictive distribution (in practice it conditions on the current masked sequence $\tilde{\mathbf{a}}_{t}$ and $\mathbf{c}$). Thus, at each step, Discrete Diffusion VLA recovers only a subset of masked positions and leaves the rest masked, moving from higher to lower mask ratios until reaching $\mathbf{a}_{0}$.
@@
In implementation, we follow masked-diffusion / infilling formulations and collapse the multi-step chain into a single masked-token prediction objective. We draw a mask ratio $\gamma_{t}\in(0,1]$ that emulates a time $t$, replace the selected action positions by special token [MASK] to obtain $\tilde{\mathbf{a}}_{t}$, and train a transformer $f_{\theta}$ to predict the original tokens with cross-entropy on masked indices:
@@
$\mathcal{L}_{\text{CE}}(\theta)\;=\;-\sum_{i\in\mathcal{M}_{\gamma_{t}}}\log p_{\theta}\bigl{(}a_{0,i}\mid\tilde{\mathbf{a}}_{t},\mathbf{c}\bigr{)},\qquad p_{\theta}(\cdot)=\mathrm{softmax}\bigl{(}W\,f_{\theta}(\tilde{\mathbf{a}}_{t},\mathbf{c})\bigr{)},\vskip-3.0pt$ (6)
@@
where $\mathcal{M}_{\gamma_{t}}$ is the masked set and $W$ projects hidden states of action positions to the $K$-way action vocabulary. This objective preserves diffusion’s corruption–denoising spirit while using a simple maximum-likelihood surrogate. As shown in recent analyses of masked diffusion (Shi et al., 2024a; Kim et al., 2025a), such losses upper-bound the negative log-likelihood under appropriate schedules. At inference time, we instantiate the reverse process via parallel decoding with adaptive re-masking: (i) initialize all $L$ action positions as [MASK]; (ii) predict token distributions for all positions in parallel; (iii) keep high-confidence predictions and re-mask the uncertain ones; (iv) repeat for a small number of iterations.
@@
Discrete Diffusion VLA accepts increased training‑time complexity (i.e., solving exponentially many infilling tasks) to gain arbitrary-order decoding at test time, selecting the inference order adaptively (e.g., by confidence or confidence gap), different from BERT-style masked LM which uses a fixed, small mask ratio in a single pass and lacks a principled generative reverse chain.
3.3 Unified VLA Architecture
The architecture of our Discrete Diffusion VLA is shown in Fig. 2.

### v2

@@ Abstract
Vision–Language–Action (VLA) models adapt large vision–language backbones to map images and instructions into robot actions. However, prevailing VLAs either generate actions autoregressively in a fixed left-to-right order or attach separate MLP or diffusion heads outside the backbone, leading to fragmented information pathways and specialized training requirements that hinder a unified, scalable architecture. We present Discrete Diffusion VLA, a unified-transformer policy that models discretized action chunks with discrete diffusion. The design retains diffusion’s progressive refinement paradigm while remaining natively compatible with the discrete token interface of VLMs. Our method achieves an adaptive decoding order that resolves easy action elements before harder ones and uses secondary re-masking to revisit uncertain predictions across refinement rounds, which improves consistency and enables robust error correction. This unified decoder preserves pretrained vision-language priors, supports parallel decoding, breaks the autoregressive bottleneck, and reduces the number of function evaluations. Discrete Diffusion VLA achieves 96.3% avg. success rates on LIBERO, 71.2% visual matching on SimplerEnv-Fractal and 54.2% overall on SimplerEnv–Bridge, improving over autoregressive, MLP decoder and continuous diffusion baselines. These findings indicate that discrete-diffusion VLA supports precise action modeling and consistent training, laying groundwork for scaling VLA to larger models and datasets. Our code will be available at https://github.com/Liang-ZX/DiscreteDiffusionVLA.
@@ 1 Introduction
[…]
@@
Drawing on recent advances of discrete diffusion and discrete flow‑matching for language and multi-modal generation (Nie et al., 2025a; Shi et al., 2024b; Gat et al., 2024; Kim et al., 2025a; Yang et al., 2025a), we introduce Discrete Diffusion VLA, the first VLA framework to unify vision, language, and action in a single transformer, maintaining strong VLM priors and achieving precise action modeling. In Discrete Diffusion VLA, each action dimension is first discretized into tokens via binning scheme and then grouped into fixed-length chunks. The fixed-length token generation is exactly suitable for discrete diffusion models. During training, we mask a subset of tokens in the action chunk and train the transformer to predict them from the context of unmasked tokens across all modalities. At inference, we start with all action tokens masked and iteratively predict them and re-mask low-confident ones until convergence, in line with the “first‑easy, then‑hard” philosophy. Moreover, we employ a secondary re-masking technique to guarantee the consistency across different denoising steps, yielding flexible parallel decoding and robust error correction.
@@
For robotics manipulation tasks, our Discrete Diffusion VLA keeps action generation inside a unified transformer with the same training objective (i.e., cross-entropy loss) as VLM, different from separate action decoder methods. This preserves much of the backbone’s pretrained vision and language capability (analogous to extending a LLM to new languages), while offering a potential path to inherit from unified transformer’s scaling behavior, paving the way for future large‑scale VLA research. On the other hand, Discrete Diffusion VLA breaks AR model’s left-to-right bottleneck. Action chunks are adaptively decoded in parallel over a small number of steps, and unconfident tokens can be revisited via iterative re-masking, leveraging full cross-modal context (including actions itself) for refinement.
@@
We evaluate Discrete Diffusion VLA across multiple robots and tasks: (1) a Franka Panda arm on LIBERO (Liu et al., 2023), (2) a Google Robot on SimplerEnv–Fractal (Li et al., 2025), and (3) a WidowX Arm on SimplerEnv–Bridge (Li et al., 2025). Using only RGB inputs, language and end-effector positions (no depth and affordances), Discrete Diffusion VLA achieves 96.3% average success rates on LIBERO ( +0.9% vs. OpenVLA-OFT (Discrete) ), 71.2% visual matching with 64.1% overall on SimplerEnv–Fractal, and 54.2% overall on SimplerEnv–Bridge (+14.7% over $\pi_{0}$ and +6.4% over $\pi_{0}$-FAST). Our model outperforms all of AR, MLP decoder and diffusion baselines across tasks while using fewer number of function evaluations (NFEs) than AR. Ablations confirm the benefits of our adaptive decoding strategy.
@@
In summary, our contributions are threefold: 1) We introduce the first discrete diffusion VLA, unifying action generation with vision–language in one transformer while maintaining strong performance. 2) We develop an adaptive decoding strategy with iterative re-masking, enabling parallel action-token decoding and error correction, improving accuracy with the unified architecture. 3) We validate Discrete Diffusion VLA on Franka Panda, Google Robot and WidowX, achieving 96.3% avg. SR on LIBERO, 64.1% and 54.2% overall on SimplerEnv–Fractal and –Bridge, consistently outperforming AR, MLP decoder and continuous diffusion baselines (e.g., $\pi_{0}$ and $\pi_{0}$-FAST).
@@ 2 Related Works
2.1 Vision-Language-Action Models
Early VLA systems took a two-part form: RT-1 and RT-2 first let the VLM produce latent tokens, then a separate MLP action decoder maps them to discretized controls in a single shot (Brohan et al., 2022; Zitkovich et al., 2023). Subsequent work shifted to AR token policies and scaled backbones and vision encoders for general manipulation (e.g. OpenVLA) (Kim et al., 2024; Touvron et al., 2023b; Oquab et al., 2024a; Zhai et al., 2023), while learning latent action shows better versatility (e.g. LAPA (Ye et al., 2024)). For high-frequency control and fast adaptation, token compression and action chunking are introduced and improve efficiency (e.g. $\pi_{0}$-FAST and OpenVLA-OFT) (Pertsch et al., 2025; Kim et al., 2025b; Song et al., 2025). In parallel, diffusion and flow matching action heads model continuous trajectories (Janner et al., 2022; Chi et al., 2023; Liang et al., 2023; Liu et al., 2024; Ni et al., 2023; Liang et al., 2025; Shen et al., 2025), spanning lightweight transformers (Li et al., 2024a), and hierarchical designs (e.g. $\pi_{0}$$\pi_{0.5}$(Black et al., 2024; Intelligence et al., 2025; Bu et al., 2025b; a; Wen et al., 2025; Zhong et al., 2025; Liu et al., 2025; Liang et al., 2024; Yang et al., 2025b), typically via a separate denoising loop conditioned on language and vision tokens.
In contrast, we perform discrete diffusion inside a single transformer over tokenized action chunks, enabling parallel, revisable decoding via iterative re-masking in a few steps while preserving the VLM’s emergent multimodal capabilities.
@@ 2.2 Discrete Diffusion Models
Discrete diffusion has recently achieved strong results on discrete data, notably tokenized images and natural language. Foundational work like D3PM (Austin et al., 2021) formalizes discrete diffusion as a Markov chain. Different from continuous diffusion models, it factorizes across positions, so each token (represented as a one-hot vector) is independently corrupted into a categorical distribution. VQ-Diffusion (Gu et al., 2022) and MaskGIT (Chang et al., 2022a) build on this, achieving high-fidelity image generation with transformers that iteratively predict masked/corrupted image tokens in a non-autoregressive manner. In language, Diffusion-BERT (He et al., 2022) and subsequent Masked Diffusion Models (Shi et al., 2024c; Zheng et al., 2024) demonstrate the viability of this approach and more recent LLaDA (Nie et al., 2025b) and DiffuLLaMA (Gong et al., 2024) scale it to 7B LMs competitive with AR baselines. MMaDA (Yang et al., 2025a) further develops this direction and shows the unified discrete diffusion model can jointly generate text and images. Concurrently, VPDD (He et al., 2024) explores discrete diffusion for policy learning via large-scale video-prediction pretraining and action fine-tuning, but it does not instantiate a unified VLA.
Our work extends this line to the VLA model: we perform discrete diffusion over tokenized action chunks, preserving language capabilities and VLM synergy, yielding competitive state-of-the-art VLA performance and laying groundwork toward unified, scalable, multi-modal foundation models spanning vision, language, and actions.
@@ 3.1 Overview
Figure 2 outlines our Discrete Diffusion VLA. We cast action decoding as discrete diffusion via masked-token denoising inside the same transformer that encodes vision and language. Given image observations (single- or multi-view) and a language instruction, each continuous control dimension is discretized into tokens and packed into a fixed-length future action chunk. A single transformer attends to frozen visual features and pretrained-LM text embeddings while progressively unmasking action tokens according to a diffusion schedule, so perception, instruction grounding, and action denoising are executed within one unified model.
@@ 3.2 Discrete Diffusion over Action Tokens
Let a single action chunk be a length-$L$ token sequence $\mathbf{a}_{0}=(a_{0,1},\dots,a_{0,L})$, where each $a_{0,i}\in\{1,\dots,K\}$ is obtained by binning continuous controls (position, orientation, etc.) following prior methods (Brohan et al., 2022; Kim et al., 2024). We augment the action vocabulary with a special mask token M (i.e., [MASK]), yielding V=K+1V=K{+}1 symbols and one-hot basis $\{\mathbf{e}_{1},\dots,\mathbf{e}_{K},\mathbf{e}_{\textrm{M}}\}$.
@@
Composing transition matrices yields $\bar{\mathbf{Q}}_{t}=\mathbf{Q}_{t}\cdots\mathbf{Q}_{1}$, and the corrupted distribution at time $t$ factorizes across positions with $L$ the length of action chunk tokens.
@@
The reverse (denoising) process defines conditionals $p_{\theta}(\mathbf{a}_{t-1}\mid\mathbf{a}_{t},\mathbf{c})$ under multimodal (i.e., vision and language) context $\mathbf{c}$. By Bayes’ rule, for each position $i$, we have
@@
where $\pi_{\theta}(i\mid\mathbf{c})$ is the model’s predictive distribution. Thus, at each step, Discrete Diffusion VLA recovers only a subset of masked positions and leaves the rest masked, moving from higher to lower mask ratios until reaching $\mathbf{a}_{0}$.
@@
In implementation, we follow mask diffusion formulations and collapse the multi-step chain into a single masked-token prediction objective. We draw a mask ratio $\gamma_{t}\in(0,1]$ that emulates diffusion time $t$, replace the selected action positions by special token [MASK] to obtain $\tilde{\mathbf{a}}_{t}$, and train a transformer $f_{\theta}$ to predict the original tokens with cross-entropy on masked indices:
@@
$\mathcal{L}_{\text{CE}}(\theta)\;=\;-\sum_{i\in\mathcal{M}_{\gamma_{t}}}\log p_{\theta}\bigl(a_{0,i}\mid\tilde{\mathbf{a}}_{t},\mathbf{c}\bigr),\qquad p_{\theta}(\cdot)=\mathrm{softmax}\bigl(W\,f_{\theta}(\tilde{\mathbf{a}}_{t},\mathbf{c})\bigr),$ (5)
@@
where $\mathcal{M}_{\gamma_{t}}$ is the masked set and $W$ projects hidden states of action positions to the $K$-way action local vocabulary. This objective preserves diffusion’s corruption–denoising spirit while using a simple maximum-likelihood surrogate. As shown in recent analyses (Shi et al., 2024a; Kim et al., 2025a), such losses upper-bound the negative log-likelihood under appropriate schedules.
@@
Discrete Diffusion VLA accepts increased training‑time complexity (i.e. solving exponentially many infilling tasks) to gain arbitrary-order decoding at test time, selecting the inference order adaptively by confidence or confidence gap, different from BERT-style parallel decoding which uses a fixed, small mask ratio in a single pass and lacks a principled generative reverse chain.
3.3 Our Unified VLA Architecture